OpenStudy (jakeenglish):

how do you get 2i to equal zero? i'm writing a polynomial function, and 2i is one of the given zeros.

OpenStudy (kainui):

So specifically you want your polynomial to be zero when you plug in x=2i is that correct? Then all you have to do is take it, and change it: \[x=2i\] subtract 2i from both sides, \[x-2i=0\] You could multiply other things by this if you want, but now you can see that: \[f(x)=x-2i\] so now you have: \[f(2i)=0\]